OpenStudy (anonymous):

please help if y=4x^3-2x determine dy/dt when x=-2 and dx/dt=-2

OpenStudy (anonymous):

@jim_thompson5910 @freckles

OpenStudy (freckles):

\[\frac{dy}{dx}=\frac{\frac{dy}{dt}}{\frac{dx}{dt}}\]

OpenStudy (freckles):

\[\frac{dy}{dx} \cdot \frac{dx}{dt}=\frac{dy}{dt}\]

OpenStudy (freckles):

since you wanted to solve for dy/dt anyways

OpenStudy (freckles):

you are given dx/dt

OpenStudy (freckles):

find dy/dx by differentiating y=4x^3-2x w.r.t x

OpenStudy (anonymous):

i got -92

OpenStudy (anonymous):

thanks